OpenStudy (anonymous):

Sally was asked to find two integers such that one is three less than 4 times the other and the sum of the two integers is 42. The integers must be _____. A. 14, 28 B. 15, 27 C. 9, 33 D. -13, 55

OpenStudy (studygurl14):

Use these equations: \[i_{1}\] means the first integer \[i_{2}\] means the second integer \[i_{1} = 4i _{2} - 3\] \[i_{1} + i_{2} = 42\]

OpenStudy (anonymous):

can u help me

OpenStudy (studygurl14):

sure. with what?

OpenStudy (anonymous):

im just confuse i had try to get the answer with the equations that u gave me but i dont get the answer

OpenStudy (anonymous):

i try

OpenStudy (studygurl14):

okay, so do this \[(4(i_{2}) - 3)+ i_{2} = 42\]

OpenStudy (anonymous):

4(-13) - 3 =-55

OpenStudy (anonymous):

the second integer is what the second # of the multiple choice

OpenStudy (studygurl14):

I'm confused about what you're doing...

OpenStudy (anonymous):

yeah every body is even i can u plz tel me the answer plz in my question i gave u the mutiple choices plz

OpenStudy (texaschic101):

x = 1st integer y = 2nd integer x = 4y - 3 x + y = 42 4y - 3 + y = 42 5y - 3 = 42 5y = 42 + 3 5y = 45 y = 9 x = 4y - 3 x = 4(9) - 3 x = 36 - 3 x = 33 check.. x = 4y - 3 33 = 4(9) - 3 33 = 36 - 33 33 = 33 (correct) integers are 9 and 33